# Magnetic reversals in a geodynamo model with a stably-stratified layer

### v1

@@ 1 Introduction
Planetary magnetic fields are generated and maintained through the geodynamo process, mechanism through which kinetic energy is converted into magnetic energy, being thus able to sustain a large scale magnetic field (Moffatt, 1978; Landeau et al., 2022; Jones, 2011). This mechanism takes place in the core of planets, where the turbulent motion of an electrically conducting fluid is mostly driven by convection and strongly affected by rotation. In the solar system, most planets produce a self-sustained magnetic field, such as the Earth, Mercury and the gaseous giants. Due to differences in the internal composition of these planets, the properties of their magnetic fields vary, including their intensity, tilt angle relative to the rotation axis, degree of axial symmetry, and north-south symmetry, among others. The Earth has a magnetic field with a strong dipolar component with tilting angle of about $11^{\circ}$ and an intensity at its surface that varies between $0.25$ and $0.65$ G.
@@
One of the most interesting phenomena of the geomagnetic field is its ability to reverse its polarity in an irregular and chaotic manner (Lowrie and Kent, 2013). Over the past $80$ million years, Earth’s magnetic field has reversed on average with a frequency of approximately $4$ Myr-1, with individual reversal events typically lasting around $20,000$ years. However, both the frequency and duration of reversals vary significantly: some intervals, known as superchrons, can span tens of millions of years without any reversals, while others show more frequent and irregular polarity changes. Understanding this process has been one of the main objectives of direct numerical simulations (DNS) of the geodynamo model (Braginsky and Roberts, 1995; Glatzmaier and Roberts, 1995; Christensen et al., 2001; Kutzner and Christensen, 2002; Li et al., 2002; Wicht and Olson, 2004; Takahashi et al., 2005; Driscoll and Olson, 2009; Olson et al., 2009; Amit et al., 2010; Olson et al., 2011; Olson and Amit, 2014; Sheyko et al., 2016; Menu et al., 2020). Parametric studies on this problem suggested that the transition from stable dynamos dominated by a strong dipolar component to reversing solutions is controlled by the turbulent strength of the flow (Kutzner and Christensen, 2002; Christensen and Aubert, 2006). As the flow becomes more turbulent, the reversal frequency increases. More precisely, this transition has been proposed to be controlled by the local Rossby number defined as $\mathrm{Ro}_{\ell}=\mathrm{Ro}\bar{\ell}_{\mathrm{u}}/\pi$, with $\mathrm{Ro}=U/(L\Omega)$ the Rossby number, with $U$ the characteristic velocity, $L$ the characteristic length scale, $\Omega$ the rotation rate, and $\bar{\ell}_{\mathrm{u}}$ the characteristic spherical harmonic degree of the flow. Numerical simulations first suggested that the transition takes place around a critical value $\mathrm{Ro}_{\ell}\approx 0.1$, with smaller values of $\mathrm{Ro}_{\ell}$ corresponding to stable dipolar solutions. The reversals found for values $\mathrm{Ro}_{\ell}>0.1$ are typically non-dipolar, meaning that the dipole component is relatively small with respect to the quadrupole and higher order modes, which make these kind of reversals not Earth-like. This regime is usually known as multipolar dynamo. There has been since then an interest to find a regime of parameters in which Earth-like magnetic reversals are observed. In particular, simulations show that Earth-like magnetic field reversals occur only within a very narrow range of local Rossby number around $0.1$, a degree of fine-tuning that seems unlikely to have been maintained over hundreds of millions of years in Earth’s core. Furthermore, it was later observed that the local Rossby number fails to describe the dipolar-multipolar transition for some range of parameters, and other control parameters have been proposed, such as the ratio between kinetic and magnetic energies (Menu et al., 2020; Tassin et al., 2021; Zaire et al., 2022; Frasson et al., 2025).
@@
The exact nature of convective motion in Earth’s outer core remains uncertain. Standard models typically assume that convection occurs throughout the entire region between the inner-core boundary (ICB) and the outer-core boundary (CMB). However, recent seismic and geomagnetic measurements have questioned this assumption, suggesting the existence of a stably-stratified layer (SSL) below the CMB (Helffrich and Kaneshima, 2010, 2013; Buffett et al., 2016; Kaneshima, 2018). The origin of this layer may be either thermal or compositional in nature, potentially resulting from the accumulation of light elements below the CMB or the high thermal conductivity of the liquid metal relative to the heat flux at the CMB, leading to a sub-adiabatic temperature gradient (Pozzo et al., 2012). The existence of such layer in Earth’s core remains a subject of debate, with its estimated size ranging from $\mathcal{H}_{s}=0$ km to $\mathcal{H}_{s}=450$ km, and stratification strengths ranging from $N=0$ to $N=50\Omega$, where $N$ is the buoyancy frequency (Christensen, 2018; Gastine et al., 2020). The presence of this layer has also been predicted in other planetary systems, such as Mercury, Saturn or Jupiter, often used to explain different large-scale structures of the magnetic field (Christensen, 2006; Moore et al., 2018; Takahashi et al., 2019; Yan and Stanley, 2021; Guervilly, 2022). One of the main effects of a SSL below the CMB is that high-order modes of the magnetic field are smoothed out due to the skin effect, leading to a magnetic field dominated by low-order modes.
@@
In addition to the possible presence of a SSL below the CMB, seismic measurements and laboratory experiments suggest that the heat flux at the CMB is not uniform, but it rather exhibits a complex spatial distribution (Su et al., 1994; Sumita and Olson, 2002). DNS of the geodynamo problem have shown that an heterogeneous heat flux at the CMB can modify the morphology of the magnetic field, trigger polarity reversals, and control their frequency (Glatzmaier et al., 1999; Olson et al., 2010; Gissinger et al., 2012; Christensen, 2018; Frasson et al., 2024; Terra-Nova and Amit, 2024; Yan et al., 2023). In particular, the non-axisymmetric structure of the geomagnetic field at the CMB are often associated to lateral heterogeneous heat flux.
@@ 2 Model
We model the dynamics of an electrically conducting fluid using the incompressible magnetohydrodynamics (MHD) equations under the Boussinesq approximation in a spherical shell of inner radius $r_{i}$ and outer radius $r_{o}$. The aspect ratio is defined as $\chi=r_{i}/r_{o}$ and is fixed to $0.35$ as the case of Earth. We express the length in units of the gap $L=r_{o}-r_{i}$, and the time in units of $L^{2}/\nu$, with $\nu$ the kinematic viscosity. The magnetic field $\bm{B}$ is expressed in units of $\sqrt{\rho\mu\eta\Omega}$, with $\rho$ the mean fluid density, $\mu$ the magnetic permeability, $\eta$ the magnetic diffusivity and $\Omega$ the rotation rate. The velocity field $\bm{u}$ is expressed in units of $\nu/L$ and the temperature $T$ in units of $q_{o}L/k$ with $q_{o}$ the mean heat flux at the CMB and $k$ the thermal conductivity. These units are used to obtained the dimensionless MHD equations (Braginsky and Roberts, 1995)
@@
We solve this system of equations using the PARODY-JA code (Dormy et al., 1998; Aubert et al., 2008), that uses a pseudo-spectral method in the poloidal and azimuthal directions, and a second-order finite-differences scheme for the radial direction, with a refined number of collocation points near the boundaries. The code implements a MPI parallelisation in the radial direction, and it is coupled to the open-source library SHTns (Schaeffer, 2013) that implements an openMP parallelisation in the angular directions to improve the performance of the spherical harmonic transforms. The resolution in the angular directions is defined such that the ratio between the maximum and minimum values of the magnetic energy spectrum is larger than $50$. The resolution in the radial direction is chosen so that there are at least $10$ points in the Ekman layer at the boundaries.
We introduce an SSL in our system of equations in an ad-hoc manner by implementing a static temperature gradient background given by a piecewise function. In the convective region, this gradient is given by the solution of $\nabla^{2}T_{s}=0$ for $r_{i}<r<r_{s}$, while we propose a simple linear profile for the stably-stratified region $r_{s}<r<r_{o}$, leading to (Nakagawa, 2011, 2015; Couston et al., 2017; Christensen, 2018; Gastine et al., 2020)
@@
with $\mathcal{H}_{s}=r_{o}-r_{s}$ the size of the SSL, and $\Gamma$ an extra free parameter that controls the maximum stratification strength as $N_{\mathrm{max}}/\Omega=\sqrt{\mathrm{Ra}\mathrm{E}\Gamma}$. We choose this value so that the maximum stratification strength is between $5$ and $15$, corresponding to a relatively strong stratification. We vary the size of the layer from $\mathcal{H}_{s}=0$ to $\mathcal{H}_{s}=0.24L$. The shape of the static temperature gradient is shown schematically in Fig. 1. Other choices of the static temperature gradient are possible, such as a continuous profile given by an hyperbolic tangent (Nakagawa, 2011, 2015; Takahashi et al., 2019). The drawback of these choices is that they introduce an extra parameter that determines the size of the transition between the convective and stably-stratified regions. This layer can also be modelled using double-diffusive convection, considering both thermal and compositional buoyancies, that leads to different regimes such as classical or finger convection (Monville et al., 2019; Guervilly, 2022). In this work, we restrict ourselves to a linear profile for the SSL below the CMB.
@@ 3 Outputs
Figure 3: Dipolar-multipolar transition for different sizes $\mathcal{H}_{s}$ of the stably-stratified layer as a function of (a) the Rayleigh number and (b) the local Rossby number for $\mathrm{E}=10^{-3}$, $\Pr=1$, and $\mathrm{Pm}=10$.
The code PARODY-JA solves the incompressible MHD Eqs. (1)-(4) using the poloidal-toroidal decomposition, where the total magnetic field $\bm{B}$ is expressed in terms of two scalar potentials $B_{\mathrm{pol}}$ and $B_{\mathrm{tor}}$ such that $\bm{B}=\bm{\nabla}\times\bm{\nabla}\times(B_{\mathrm{pol}}\bm{r})+\bm{\nabla}\times(B_{\mathrm{tor}}\bm{r})$. The total magnetic field $\bm{B}$ can be described as a linear combination of spherical harmonics $Y_{l}^{m}$. In particular, we can extract the axisymmetric dipole and quadrupole components of the total magnetic field from the poloidal and toroidal decomposition as
$\begin{split}D&=B_{\mathrm{pol}}(l=1,m=0)+B_{\mathrm{tor}}(l=2,m=0),\\ Q&=B_{\mathrm{pol}}(l=2,m=0)+B_{\mathrm{tor}}(l=1,m=0).\end{split}$ (7)
@@
with $V$ the volume of the spherical shell. The amplitudes of the velocity and magnetic fields are quantified by the magnetic Reynolds number $\mathrm{Rm}=UL/\eta$ and the Elsasser number $\Lambda=B^{2}/(\rho\mu\eta\Omega)$ with $U$ and $B$ the volume-averaged RMS amplitudes of the velocity and magnetic fields, respectively. For the adimensionalisation used in our simulations, these numbers are computed as
$\mathrm{Rm}=\mathrm{Pm}\sqrt{2E_{\mathrm{kin}}},$ (11)
$\Lambda=2\mathrm{Pm}\mathrm{E}E_{\mathrm{mag}}.$ (12)
@@
$f_{\mathrm{dip}}=\frac{E_{\mathrm{mag}}^{\mathrm{CMB}}(l=1)}{\sum_{l=1}^{12}E_{\mathrm{mag}}^{\mathrm{CMB}}(l)}$ (13)
with $E_{\mathrm{mag}}^{\mathrm{CMB}}(l)=\sum_{m}E_{\mathrm{mag}}^{\mathrm{CMB}}(l,m)$ the magnetic energy spectrum evaluated at the CMB. The denominator is truncated at $l=12$ to compare with geomagnetic observations, that lead to a typical value for the Earth $f_{\mathrm{dip}}=0.68$ (Christensen and Aubert, 2006; Menu et al., 2020). The transition is typically controlled by the local Rossby number
@@
We also measure the equatorial symmetry of the magnetic field using the mean hemisphericity at the CMB
@@
with $B_{r}^{N}$ and $B_{r}^{S}$ the radial magnetic field evaluated at the CMB in the northern and southern hemispheres, respectively. For an equatorially symmetric system, the mean hemisphericity is zero, while if the magnetic field is concentrated in one hemisphere, it is $\pm 1$.
@@ 4 Results
We perform a parametric study of the geodynamo fixing $\Pr=1$, $\mathrm{E}=10^{-3}$, $\mathrm{Pm}=10$, the aspect ratio of the system $\chi=0.35$, and varying $\mathrm{Ra}$, and the size of the SSL $\mathcal{H}_{s}=r_{o}-r_{s}$ keeping the stratification strength between $N_{\mathrm{max}}/\Omega\in(5,15)$. Some simulations are started with a fluid at rest, imposing a small temperature and magnetic perturbations, while some others are restarted from a simulation with similar parameters.
@@
The effects of the SSL in the magnetic field morphology are shown in Fig. 2. The dipolar strength of the magnetic field at the CMB $f_{\mathrm{dip}}$ defined in Eq. (13) increases with the size of the SSL for a fixed value of $\mathrm{Ra}/\mathrm{Ra}_{c}$, where $\mathrm{Ra}_{c}$ is the numerically determined critical Rayleigh number for the onset of convection. This is a consequence of the skin effect, which attenuates the contribution of fast modes of the magnetic field. The visualisation of the radial velocity and magnetic fields in Fig. (b)-(c) show how they are attenuated in the presence of the SSL.

### v3

@@ 1 Introduction
Planetary magnetic fields are generated and maintained through the dynamo process, mechanism through which kinetic energy is converted into magnetic energy, being thus able to sustain a large scale magnetic field (Moffatt1978; Landeau2022; Jones2011). This mechanism takes place in the core of planets, where the turbulent motion of an electrically conducting fluid is mostly driven by convection and strongly affected by rotation. In the solar system, most planets produce a self-sustained magnetic field, such as the Earth, Mercury and the gaseous giants. Due to differences in the internal composition and size of these planets, the properties of their magnetic fields vary, including their intensity, tilt angle relative to the rotation axis, degree of axial symmetry, and north-south symmetry, among others. At present, the Earth has a magnetic field with a strong dipolar component with tilting angle of about $11^{\circ}$ and an intensity at its surface that varies between $0.25$ and $0.65$ G.
@@
One of the most interesting phenomena of the geomagnetic field is its ability to reverse polarity in an irregular and chaotic manner (Lowrie2013). Over the past $80$ million years, Earth’s magnetic field has reversed on average with a frequency of approximately $4$ Myr-1, with individual reversal events typically lasting around $20,000$ years. However, both the frequency and duration of reversals vary significantly: some intervals, known as superchrons, can span tens of millions of years without any reversals, while others show more frequent and irregular polarity changes. Understanding this process has been one of the main objectives of direct numerical simulations (DNS) of the geodynamo model (Braginsky1995; Glatzmaier1995; Christensen2001; Kutzner2002; Li2002; Wicht2004; Takahashi2005; Driscoll2009; Olson2009; Amit2010; Olson2011; Olson2014; Sheyko2016; Menu2020). Parametric studies on this problem suggested that the transition from stable dynamos dominated by a strong dipolar component to reversing solutions is controlled by the relative amplitude of inertia and Coriolis force (Kutzner2002; Christensen2006). As the flow becomes more turbulent, the reversal frequency increases. More precisely, this transition has been proposed to be controlled by the local Rossby number defined as $\mathrm{Ro}_{\ell}=\mathrm{Ro}\bar{\ell}_{\mathrm{u}}/\pi$, with $\mathrm{Ro}=U/(L\Omega)$ the Rossby number, with $U$ the characteristic velocity, $L$ the characteristic length scale, $\Omega$ the rotation rate, and $\bar{\ell}_{\mathrm{u}}$ the characteristic spherical harmonic degree of the flow. Numerical simulations first suggested that the transition takes place around a critical value $\mathrm{Ro}_{\ell}\approx 0.1$, with smaller values of $\mathrm{Ro}_{\ell}$ corresponding to stable dipolar solutions. The reversals found for values $\mathrm{Ro}_{\ell}>0.1$ are typically non-dipolar, meaning that the dipole component is relatively small with respect to the quadrupole and higher order modes, which make these kind of reversals not Earth-like. This regime is usually known as multipolar dynamo. There has been since then an interest to find a regime of parameters in which Earth-like magnetic reversals are observed. In particular, simulations show that Earth-like magnetic field reversals occur only within a very narrow range of local Rossby number around $0.1$, a degree of fine-tuning that seems unlikely to have been maintained over hundreds of millions of years in Earth’s core. Furthermore, it was later observed that the local Rossby number fails to describe the dipolar-multipolar transition for some range of parameters, and other control parameters have been proposed, such as the ratio between kinetic and magnetic energies (Menu2020; Zaire2022; Frasson2025). In particular, Tassin2021 showed that reversing dynamos occur when this ratio is larger than $\mathcal{O}(1)$, which is much larger than the typical values estimated for the Earth’s core of $10^{-3}$. In the recent work of Jones2025, it was shown that it is possible to obtain Earth-like reversals while maintaining realistic kinetic-to-magnetic energy ratios by increasing the magnetic Prandtl number together with a corresponding increase in the thermal Prandtl number, preserving realistic transport levels.
@@
The exact nature of convective motion in Earth’s outer core remains uncertain. Standard models typically assume that convection occurs throughout the entire region between the inner-core boundary (ICB) and the core-mantle boundary (CMB). However, recent seismic and geomagnetic measurements have questioned this assumption, suggesting the existence of a stably-stratified layer (SSL) below the CMB (Helffrich2010; Helffrich2013; Buffett2016; Kaneshima2018). The origin of this layer may be either thermal or compositional in nature, potentially resulting from the accumulation of light elements below the CMB or the high thermal conductivity of the liquid metal relative to the heat flux at the CMB, leading to a sub-adiabatic temperature gradient (Pozzo2012). The existence of such layer in Earth’s core remains a subject of debate, with its estimated size ranging from $\mathcal{H}_{s}=0$ km to $\mathcal{H}_{s}=450$ km, and stratification strengths ranging from $N=0$ to $N=50\Omega$, where $N$ is the buoyancy frequency (Christensen2018; Gastine2020). The presence of this layer has also been predicted in other planetary systems, such as Mercury, Saturn or Jupiter, often used to explain different large-scale structures of the magnetic field (Christensen2006a; Moore2018; Takahashi2019; Yan2021; Guervilly2022). One of the main effects of a SSL below the CMB is that high-order modes of the magnetic field are smoothed out due to the skin effect, leading to a magnetic field dominated by low-order modes. The recent work of Aubert2025 proposed a mechanism that differs from the classical approaches based on the levels of inertia of the flow. The proposed kinematic mechanism suggests that in the presence of a stable top layer, there is a competition between the subsurface upwelling (generated in the convective region) and the surface circulation (generated in the stable top layer).
@@
In addition to the possible presence of a SSL below the CMB, seismic measurements and laboratory experiments suggest that the heat flux at the CMB is not uniform, but it rather exhibits a complex spatial distribution (Su1994; Sumita2002). DNS of the geodynamo problem have shown that an heterogeneous heat flux at the CMB can modify the morphology of the magnetic field, trigger polarity reversals, and control their frequency (Glatzmaier1999; Olson2010; Gissinger2012; Christensen2018; Yan2023; Terra-Nova2024; Frasson2025). In particular, the non-axisymmetric structure of the geomagnetic field at the CMB are often associated to lateral heterogeneous heat flux.
@@ 2 Model
We model the dynamics of an electrically conducting fluid using the incompressible magnetohydrodynamics (MHD) equations under the Boussinesq approximation in a spherical shell of inner radius $r_{i}$ and outer radius $r_{o}$. The aspect ratio is defined as $\chi=r_{i}/r_{o}$ and is fixed to $0.35$ as the case of Earth. We express the length in units of the gap $L=r_{o}-r_{i}$, and the time in units of $L^{2}/\nu$, with $\nu$ the kinematic viscosity. The magnetic field $\bm{B}$ is expressed in units of $\sqrt{\rho\mu\eta\Omega}$, with $\rho$ the mean fluid density, $\mu$ the magnetic permeability, $\eta$ the magnetic diffusivity and $\Omega$ the rotation rate. The velocity field $\bm{u}$ is expressed in units of $\nu/L$ and the temperature $T$ in units of $q_{o}L/k$ with $q_{o}$ the mean heat flux at the CMB and $k$ the thermal conductivity. These units are used to obtained the dimensionless MHD equations (Braginsky1995)
@@
We solve this system of equations using the PARODY-JA code (Dormy1998; Aubert2008), that uses a pseudo-spectral method in the poloidal and azimuthal directions, and a second-order finite-differences scheme for the radial direction, with a refined number of collocation points near the boundaries. The code implements a MPI parallelisation in the radial direction, and it is coupled to the open-source library SHTns (Schaeffer2013) that implements an openMP parallelisation in the angular directions to improve the performance of the spherical harmonic transforms. The resolution in the angular directions is defined such that the ratio between the maximum and minimum values of the magnetic energy spectrum is larger than $50$. The resolution in the radial direction is chosen so that there are at least $10$ points in the Ekman layer at the boundaries.
We introduce an SSL in our system of equations in an ad-hoc manner by implementing a static temperature gradient background given by a piecewise function. In the convective region, this gradient is given by the solution of $\nabla^{2}T_{s}=0$ for $r_{i}<r<r_{s}$, while we propose a simple linear profile for the stably-stratified region $r_{s}<r<r_{o}$, leading to (Nakagawa2011; Nakagawa2015; Couston2017; Christensen2018; Gastine2020)
@@
with $\mathcal{H}_{s}=r_{o}-r_{s}$ the size of the SSL, and $\Gamma$ an extra free parameter that controls the maximum stratification strength as $N_{\mathrm{max}}/\Omega=\sqrt{\mathrm{Ra}\mathrm{E}\Gamma}$. We choose this value so that the maximum stratification strength is between $5$ and $15$, corresponding to a relatively strong stratification. We vary the size of the layer from $\mathcal{H}_{s}=0$ to $\mathcal{H}_{s}=0.24L$. The shape of the static temperature gradient is shown schematically in Fig. 1. Other choices of the static temperature gradient are possible, such as a continuous profile given by an hyperbolic tangent (Nakagawa2011; Nakagawa2015; Takahashi2019). The drawback of these choices is that they introduce an extra parameter that determines the size of the transition between the convective and stably-stratified regions. This layer can also be modelled using double-diffusive convection, considering both thermal and compositional buoyancies, that leads to different regimes such as classical or finger convection (Monville2019; Guervilly2022). In this work, we restrict ourselves to a linear profile for the SSL below the CMB.
@@ 3 Outputs
Figure 3: Dipolar-multipolar transition for different sizes $\mathcal{H}_{s}$ of the stably-stratified layer as a function of (a) the Rayleigh number and (b) the local Rossby number for $\mathrm{E}=10^{-3}$, $\Pr=1$, and $\mathrm{Pm}=10$. Circles indicate stable solutions, while square markers indicate reversing solutions.
The code PARODY-JA solves the incompressible MHD Eqs. (1)-(4) using the poloidal-toroidal decomposition, where the total magnetic field $\bm{B}$ is expressed in terms of two scalar potentials $B_{\mathrm{pol}}$ and $B_{\mathrm{tor}}$ such that $\bm{B}=\bm{\nabla}\times\bm{\nabla}\times(B_{\mathrm{pol}}\bm{r})+\bm{\nabla}\times(B_{\mathrm{tor}}\bm{r})$. The total magnetic field $\bm{B}$ can be described as a linear combination of spherical harmonics $Y_{l}^{m}$. In particular, we can extract the axisymmetric dipole and quadrupole components of the magnetic field at the CMB from the poloidal and toroidal decomposition as
$\begin{split}D&=B_{\mathrm{pol}}^{\mathrm{CMB}}(l=1,m=0),\\ Q&=B_{\mathrm{pol}}^{\mathrm{CMB}}(l=2,m=0).\end{split}$ (7)
Note that the toroidal component of the magnetic field is zero at the CMB due to the insulating mantle.
@@
with $V$ the volume of the spherical shell. The amplitudes of the velocity and magnetic fields are quantified by the Reynolds number $\mathrm{Re}=UL/\nu$ and the Elsasser number $\Lambda=B^{2}/(\rho\mu\eta\Omega)$ with $U$ and $B$ the volume-averaged RMS amplitudes of the velocity and magnetic fields, respectively. For the adimensionalisation used in our simulations, these numbers are computed as
$\mathrm{Re}=\sqrt{2E_{\mathrm{kin}}},$ (11)
$\Lambda=2\mathrm{Pm}\mathrm{E}E_{\mathrm{mag}},$ (12)
from which we can compute the magnetic Reynolds number as
$\mathrm{Rm}=\mathrm{Pm}\mathrm{Re}.$ (13)
@@
$f_{\mathrm{dip}}=\frac{\sqrt{E_{\mathrm{mag}}^{\mathrm{CMB}}(l=1)}}{\sqrt{\sum_{l=1}^{12}E_{\mathrm{mag}}^{\mathrm{CMB}}(l)}}$ (14)
with $E_{\mathrm{mag}}^{\mathrm{CMB}}(l)=\sum_{m}E_{\mathrm{mag}}^{\mathrm{CMB}}(l,m)$ the magnetic energy spectrum evaluated at the CMB. The denominator is truncated at $l=12$ to compare with geomagnetic observations, that lead to a typical value for the Earth $f_{\mathrm{dip}}=0.68$ (Christensen2006; Menu2020). The transition is typically controlled by the local Rossby number
@@
To quantify the equatorial symmetry of the magnetic field, we use the mean hemisphericity at the CMB
@@
with $B_{r}^{N}$ and $B_{r}^{S}$ the radial magnetic field evaluated at the CMB in the northern and southern hemispheres, respectively. For an equatorially symmetric system, the mean hemisphericity is zero, while if the magnetic field is concentrated in one hemisphere, it is $\pm 1$. We quantify the symmetry breaking for the flow with respect to the equator from the antisymmetric toroidal and poloidal components of the kinetic energy as
$E_{\mathrm{kin}}^{\mathrm{asym}}=\sum_{l+m=2n}E_{\mathrm{kin}}^{\mathrm{tor}}(l,m)+\sum_{l+m=2n+1}E_{\mathrm{kin}}^{\mathrm{pol}}(l,m).$ (17)
@@ 4.1 Parametric study
We perform a parametric study of the geodynamo by fixing $\Pr=1$, $\mathrm{E}=10^{-3}$, $\mathrm{Pm}=10$, the aspect ratio of the system $\chi=0.35$, and varying $\mathrm{Ra}$, and the size of the SSL $\mathcal{H}_{s}=r_{o}-r_{s}$ keeping the stratification strength between $N_{\mathrm{max}}/\Omega\in(5,15)$. Some simulations are initialised with a fluid at rest, imposing a small temperature and magnetic perturbations, while others are restarted from runs with similar parameters. The relative large values of $\mathrm{E}$ and $\mathrm{Pm}$ are chosen to enable an extensive parametric study of the system, allowing us to identify different regimes of geodynamo behaviour. The input and ouptut quantities of the simulations are shown in Table 1.
Table 1: Table of simulations homogeneous boundary conditions. We report the size of the stably-stratified layer $\mathcal{H}_{s}$, the normalised Rayleigh number $\mathrm{Ra}/\mathrm{Ra}_{c}$, the kinetic energy $E_{\mathrm{kin}}$, the magnetic energy $E_{\mathrm{mag}}$, the ratio of kinetic and magnetic energies $E_{\mathrm{kin}}/E_{\mathrm{mag}}$, the magnetic Reynolds number $\mathrm{Rm}$, the antisymmetric kinetic energy $E_{k}^{\mathrm{asym}}$, the dipolar strength $f_{\mathrm{dip}}$, whether there are reversals, and the final time of the simulations $t_{\mathrm{run}}$ expressed in diffusive time units. All simulations correspond to $\mathrm{E}=10^{-3}$, $\Pr=1$, $\mathrm{Pm}=10$, and $r_{i}/r_{o}=0.35$.
$\mathcal{H}_{s}/L$ $\mathrm{Ra}/\mathrm{Ra}_{c}$ $\Lambda$ $E_{\mathrm{kin}}$ $E_{\mathrm{mag}}$ $E_{\mathrm{kin}}/E_{\mathrm{mag}}$ $\mathrm{Rm}$ $\mathrm{Ra}_{c}$ $E_{k}^{\mathrm{asym}}$ $f_{\mathrm{dip}}$ reversals? $t_{\mathrm{run}}$
0 5 16.36 94 818 0.12 137 47.8 0.16 0.68 no 4.2
0 10 24.12 247 1206 0.20 222 47.8 0.19 0.56 no 2
0 15 23.98 428 1198 0.36 292 47.8 0.21 0.49 no 2
0 20 23.04 625 1152 0.54 353 47.8 0.22 0.44 no 2
0 22 17.08 759 853 0.89 389 47.8 0.19 0.39 no 5
0 23 15.29 819 764 1.07 404 47.8 0.18 0.36 yes 3.75
0 25 12.89 929 644 1.44 431 47.8 0.18 0.30 yes 8
0 30 12.63 1164 631 1.84 482 47.8 0.19 0.20 yes 2
0.02 20 16.90 680 845 0.81 369 47.7 0.17 0.41 no 3
0.02 25 11.59 962 579 1.66 438 47.7 0.17 0.30 yes 6
0.04 15 22.09 417 1104 0.38 289 48.0 0.17 0.59 no 4
0.04 20 17.25 677 862 0.79 368 48.0 0.16 0.52 no 3
0.04 25 11.59 995 579 1.72 446 48.0 0.15 0.41 yes 9
0.04 30 11.96 1228 598 2.05 495 48.0 0.16 0.37 yes 3
0.09 15 22.48 373 1124 0.33 273 50.22 0.18 0.80 no 4
0.09 20 16.72 599 836 0.72 346 50.22 0.16 0.77 no 2.5
0.09 25 14.05 841 702 1.20 410 50.22 0.16 0.74 no 3
0.09 30 9.68 1175 484 2.43 484 50.22 0.16 0.53 yes 6.75
0.14 15 21.35 371 1067 0.35 272 56.81 0.18 0.87 no 2.49
0.14 20 17.29 596 864 0.69 345 56.81 0.17 0.84 no 2.54
0.14 25 15.21 841 760 1.11 410 56.81 0.16 0.83 no 2.52
0.14 30 7.76 1257 387 3.24 501 56.81 0.14 0.56 yes 2.51
0.14 35 8.74 1435 437 3.28 535 56.81 0.17 0.31 yes 1.67
0.24 15 17.43 374 871 0.43 273 66.4 0.18 0.93 no 3
0.24 20 14.28 584 713 0.82 341 66.4 0.17 0.92 no 1.27
0.24 22 13.23 681 661 1.03 369 66.4 0.17 0.92 no 3
0.24 25 9.25 927 462 2.01 430 66.4 0.15 0.89 no 7.56
0.24 26 9.35 979 467 2.10 442 66.4 0.15 0.79 yes 5.28
0.24 27.5 7.32 1100 366 3.00 469 66.4 0.15 0.56 yes 3.56
0.24 30 7.47 $12$ 373 3.25 492 66.4 0.16 0.39 yes 3
0.24 35 9.75 1342 487 2.75 518 66.4 0.19 0.34 yes 2.25
0.24 40 10.72 1508 536 2.81 549 66.4 0.20 0.33 yes 1.5
Table 2: Table of simulations with heterogeneous boundary conditions. We report the same parameters as in Table 1. The heat flux pattern at the CMB is determined by the spherical harmonic $Y_{l}^{m}$, and the amplitude of the heterogeneity is given by $\delta q$ defined in Eq. (6). All simulations correspond to $\mathrm{E}=10^{-3}$, $\Pr=1$, $\mathrm{Pm}=10$, $\mathrm{Ra}=25\mathrm{Ra}_{c}$, and a SSL of size $\mathcal{H}_{s}=0.24L$ and $N_{\mathrm{max}}/\Omega\approx 10$.
$Y_{l}^{m}$ $\delta q$ (%) $\Lambda$ $E_{\mathrm{kin}}$ $E_{\mathrm{mag}}$ $E_{\mathrm{kin}}/E_{\mathrm{mag}}$ $\mathrm{Rm}$ $E_{\mathrm{kin}}^{\mathrm{asym}}$ $f_{\mathrm{dip}}$ Reversals? $t_{\mathrm{run}}$
$Y_{1}^{0}$ -3.5 10.61 888 530 1.67 420 0.17 0.76 no 3.10
$Y_{1}^{0}$ 0.0 9.59 916 480 1.91 427 0.15 0.90 no 7.56
$Y_{1}^{0}$ 0.17 9.36 932 468 1.99 431 0.14 0.82 no 3.25
$Y_{1}^{0}$ 0.35 7.77 985 389 2.53 443 0.13 0.81 no 4.45
$Y_{1}^{0}$ 1.05 9.12 939 456 2.06 432 0.14 0.82 no 3.15
$Y_{1}^{0}$ 1.75 10.71 887 536 1.66 420 0.16 0.86 no 2.00
$Y_{1}^{0}$ 2.1 11.18 876 559 1.57 418 0.17 0.81 no 4.00
$Y_{1}^{0}$ 3.5 12.23 832 612 1.36 407 0.19 0.78 no 6.00
$Y_{1}^{0}$ 8.75 12.85 828 643 1.29 406 0.24 0.56 no 1.50
$Y_{3}^{0}$ -5.3 11.20 880 560 1.57 418 0.17 0.86 yes 9.06
$Y_{3}^{0}$ 0.0 9.42 922 471 1.96 428 0.15 0.89 no 7.56
$Y_{3}^{0}$ 0.53 10.33 884 516 1.71 420 0.16 0.92 no 5.45
$Y_{3}^{0}$ 2.65 10.72 867 536 1.64 415 0.17 0.91 no 13.44
$Y_{3}^{0}$ 5.3 9.82 908 491 1.85 425 0.16 0.83 yes 29.50
$Y_{1}^{0}+Y_{3}^{0}$ 1.75 8.54 957 427 2.24 438 0.14 0.75 yes 27.65
@@
The effects of the SSL in the magnetic field morphology are shown in Fig. 2. The dipolar strength of the magnetic field at the CMB $f_{\mathrm{dip}}$ defined in Eq. (14) increases with the size of the SSL for a fixed value of $\mathrm{Ra}/\mathrm{Ra}_{c}$, where $\mathrm{Ra}_{c}$ is the numerically determined critical Rayleigh number for the onset of convection. This is a consequence of the skin effect, which attenuates the contribution of the more spatially structured modes of the magnetic field. The visualisation of the radial velocity and magnetic fields in Figs. 2 (b)-(c) show how they are attenuated in the presence of the SSL.